Consider a hypothesis testing situation in which the null hypothesis is p = 0.05 and the alternative is p > 0.5. Data is gathered from a random sample and a test statistic of 2.1 is computed. This means the p-value would have to be:

A. less than 0.01.

B. between 0.01 and 0.05

C. between 0.05 and 0.10

D. greater than 0.10

E. More information is needed in order to determine the p-value.
Answer #1

t, po.s I too p=0.05 Test is right tailed º 2 -2.1 1 P value = el 27 2.1) = = P(2.2.1) = 1-0 19821 = 0.0179 from 2-tably betw